The probability of a defective unit in a manufacturing process has a binomial distribution with p = 0.03 What is the probability of 2 or more occurences in a sample of 15 units?
Answer #1

Let X be a number of defective units among 15 units

Here, X ~ Binomial ( n = 15, p = 0.03)

probability mass function of X is,

P(X = x) = nCx px ( 1 - p )n-x

We want to find, P(X >= 2)

P(X >= 2)

= 1 - P(X < 2)

= 1 - P(X <= 1)

= 1 - [ P(X = 0) + P(X = 1) ]

= 1 - [ 15C0 * (0.03)0 * (1-0.03)15 + 15C1 *(0.03)1 * (1-0.03)14 ]

= 1 - [ 0.63325 + 0.29378 ]

= 1 - 0.92703

= 0.07297

P(X >= 2) = 0.07297

The probability is 0.07297